OpenStudy (anonymous):

the formula to convert temperatures between degrees fahrenheit and degrees celsius is C 5/9 (F-32) apply the formula to convert 0 degrees celsius to degrees fahrenheit

OpenStudy (radar):

C=0 whats F I think somethings amiss there, I know F would be 32 lol

OpenStudy (radar):

Where is you equal sign?

OpenStudy (anonymous):

never mind i got it

OpenStudy (radar):

C=(5/9)(F-32) 0=(5/9)(F-32) F=0

OpenStudy (anonymous):

0 = 5/9 (F -32) 0* 9/5 = F - 32 0 + 32 = F so 0 deg c = 32 deg f

OpenStudy (radar):

At what temperature would F and C readings be the same?

OpenStudy (anonymous):

yea ur right harkirat i got tht

OpenStudy (cornzyblack):

i think your equation needs to be C = 5/9 ( F - 32) MAKE F the subject of formula F = 9/5(C) + 32 and substitute your value for C which is 0

OpenStudy (radar):

Just for fun.lol

OpenStudy (anonymous):

they will be same at - 40

OpenStudy (radar):

Good Harkirat

OpenStudy (anonymous):

thanks radar ☺